# XtraLight-MedMamba for Classification of Neoplastic Tubular Adenomas

### v3

@@ 5 Conclusion
In this work, we introduced XtraLight-MedMamba, an extra lightweight state-space model for the classification of Neoplastic Tubular Adenomas (NPTA). The proposed architecture is a combination of ConvNext blocks for shallow feature extraction with PVM layers to efficiently capture both short- and long-range dependencies while maintaining a minimal parameter footprint. The performance was further improved by integrating SCAB module and FNOclassifier. The proposed approach achieved enhanced classification performance on a curated dataset of low-grade tubular adenomas, which outperformed transformer-based and conventional Mamba architectures with substantially fewer parameters. Grad-CAM analysis demonstrated that the network captured subtle histological patterns, including nuclear architecture and epithelial atypia, supporting both interpretability and clinical relevance, by highlighting discriminative feature that conventional methods may overlook.
@@ Acknowledgment
The authors acknowledge the South Bend Medical Foundation (SBMF) team for their support in providing access to the Neoplastic Tubular Adenomas dataset.

### v4

@@ 5 Conclusion
In this work, we introduce XtraLight-MedMamba, an ultra-lightweight state-space model for classifying Neoplastic Tubular Adenomas (NPTA). The proposed architecture combines ConvNeXt blocks for shallow feature extraction with PVM layers to efficiently capture both short- and long-range dependencies while maintaining a minimal parameter footprint. The performance was further enhanced by integrating the SCAB module to improve cross-scale feature extraction and feature propagation. Subsequently, the FNOClassifier exhibited minimal additional trainable parameters due to their fixed orthonormal classifier directions, which imposed structured geometric constraints and reduced model complexity. The proposed approach achieved superior classification performance on a curated dataset of low-grade tubular adenomas, outperforming both transformer-based and conventional Mamba architectures while using substantially fewer parameters. Grad-CAM analysis demonstrated that the network captured subtle histological patterns, including nuclear architecture and epithelial atypia, supporting both interpretability and clinical relevance, by highlighting discriminative features that conventional methods may overlook.
@@ Acknowledgment
The authors acknowledge the South Bend Medical Foundation (SBMF) team for providing access to the Neoplastic Tubular Adenomas dataset.